Session: Sep 1997 (session 3)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (mauro at Fri Sep 26 04:27:04 METDST 1997) Success
very good, clear sky.

Cambridge (Alastair Gunn at Fri Sep 26 05:08:27 METDST 1997) Success
No known problems.

Effelsberg (dgraham at Fri Sep 26 09:56:24 METDST 1997) Minor failures
1st 3c345 scan behind hills. 1845-1902 off source.

Torun (K. Borkowski at Fri Sep 26 10:19:11 METDST 1997) Minor failures
Problems with telescope control. Off source probably after 14:30 for 15 minutes and after 19:05 for some 10 minutes.

Westerbork (Tony Foley at Fri Sep 26 17:19:30 METDST 1997) Minor failures
BIG problems with tape. First 7 passes seem OK, but the tape came off the reel somehow at about 1450UT. We tried again with this tape and still had problems. Another tape (around 1614UT) gave problems with the tape slipping on the capstan, so that the system thought that it was moving, while it was not. This gave vacuum loss after a short time. We gave up at 2111UT!

Onsala (Antonis Polatidis at Sat Sep 27 13:46:30 METDST 1997) Success
No known problems

Noto (C.Trigilio at Mon Sep 29 14:19:39 METDST 1997) Success
Noto was optional in this project because of the need of some days to assure the change of configuration from cassegrain to primary focus. However, being the weather forecast good for the 26 Sep, we observed V022c and switched Cband-> S/X on 26 Sep. No problem during V022c.

Calibration (Fredrik T Rantakyro at Wed Nov 5 15:59:18 MET 1997) Minor failures
The ANTAB formatted TSYS information can be found on VLBEER in the SEP97 directory. The file is called v022c.antab and contains the latest antenna calibration information.